OpenStudy (anonymous):

A right triangle has a hypotenuse of length 15 and a leg of length 8. What is the area of the triangle? If necessary, round your answer to two decimal places.

OpenStudy (anonymous):

http://openstudy.com/users/terenzreignz

OpenStudy (anonymous):

A right triangle has a hypotenuse of length 15 and a leg of length 8. What is the area of the triangle? If necessary, round your answer to two decimal places Hypotenuse = 15 One side =8 Let other side be x Pythagoras theorem = hyp^= side1 ^2 + side2 ^ 15^2= 8^+x^2 15^-8^2= x^2 225-64=x^2 Sqrt 161 =x Area of right triangle = ½ * side 1* side2 =½ *8*12.69 = 50.76 sq. units